OpenStudy (anonymous):

Help thanks.A survey of 100 households that own televisions found that 95 use a remote control. If you randomly knocked on two doors (from the sample of 100 households), what is the probability that the first homeowner would use a remote control and the second homeowner would not?

OpenStudy (anonymous):

Prob of first using remote = 95/100 Prob of secnd not using = 5/99 Overall prob = (95/100)(5/99) 95*5/100*99 =19/396

OpenStudy (anonymous):

thank you